In each of 4 races, the Democrats have 60% chances of winning. Assuming that the races are independent of each other, find the probability that: (1) the Democrats will win 0 races, 1 race, 2 race, 3 race, or all 4 races? (2) the Democrats will win a majority of the races.

1
Expert's answer
2022-02-01T16:52:17-0500

Let "X=" the number of wins of the Democrats out of 4 races:

X"\sim Bin(n,p)"

Given "n=4, p=0.60, q=1-p=0.40."

(1)


"P(X=0)=\dbinom{4}{0}(0.6)^0(0.4)^{4-0}=0.0256""P(X=1)=\dbinom{4}{1}(0.6)^1(0.4)^{4-1}=0.1536""P(X=2)=\dbinom{4}{2}(0.6)^2(0.4)^{4-2}=0.3456""P(X=3)=\dbinom{4}{3}(0.6)^3(0.4)^{4-3}=0.3456""P(X=4)=\dbinom{4}{4}(0.6)^4(0.4)^{4-4}=0.1296"

(2)


"P(X>2)=P(X=3)+P(X=4)""=0.3456+0.1296=0.4752"